Experimental Data 7) 1) Total Valence Electrons Rough Lewis Structure Total Valence Rough Lewis Molecular Geometry HCN actets formal charges (Indlude formal charges ather than 0) linear linear Bond Polarity (AEN) More soluble in hexane or water? Molecule Polarity Bond Polarity (AEN) More soluble in hexane or water? Molecule Polarity EN (H C)- 121-2S-04 SEN (C-N) 125-3.01 0.5 2) 8) FormulaElectrons Total Valence Rough Lewis Structure Molecular Geometry Electron Group Molecular Geometry Total Valence Electrons Formula Rough Lewis Structure OCI- widizing agent in bleach) SIF lisclude farmal charges other thon Q More soluble in hexane or water? More soluble in hexane or water? Bond Polarity [AEN) Molecule Polarity Bond Polarity (???) Molecule Polarity 9) Molecular Geometry Total Valence Rough Lewis Structure FormulaElectrons Rough Lewis Structure Electrons NoC POCI, to dissove in More soluble More soluble in hexane or water? Molecule Polarity Bond Polarity (AEN) Molecule Polarity 10) Molecular G 4) Total Valence Rough Lewis Structure FormulaElectrons Total Valence Rough Lewis Structure FormulaElectrons sOCz in chemical Molecule Polarity Molecule Polarity Bond Polarity (???)
Answer #1
TABLE 1
FORMULA TOTAL VALENCE ELECTRONS MOLECULAR GEOMETRY/ELECTRON GROUP GEOMETRY
I3- 6 because of 3 lone pairs and 2 attached atoms molecular geometry is linear and electron pair geometry is trigonal bipyramid.
OCl- 14 bent
SiF4

silicon has 4 valence electrons and fluorine has 7

total=4+7*4=32


Tetrahedral
NOCl

Nitrogen has 5 valence electron Chlorine has 7 and Oxygen has 6

total=5+7+6=18


Bent or Angular. This is due to the charge distribution on central atom.
Table 2
FORMULA BOND POLARITY MOLECULAR POLARITY MORE SOLUBLE IN WATER OR HEXANE
I3- I-I bond electronegativity difference = 2.7-2.7=0 neither polar nor non-polar because it is an ion not a molecule and ions itself are charged I3- is more soluble in water due to ion and dipole interaction between the I3- ions and water molecules.
OCl-

O-Cl bond electronegativity difference = 3.4-3.2=0.2

polar bond is present.

O---->Cl

as per like dissolve like

OCl being polar dissolves in polar only

SiF?4 bond polarity= electronegitivity of Si -electronegitivity of F=1.9-4=2.1 Though Si-F bond is polar but the SiF4 molecule isn't polar because the 4 bonds cancel the effect of each other. like dissolves like so non-polar SiF?4 dissolves in hexane.
Cl S O 3NOCl

bond polarity between N-O=electronegitivity of N-electronegitivity of O=

3-3.4=0.4

bond polarity between N-Cl=electronegitivity of N-electronegitivity of Cl=

3-3.2=0.2

NOCl is a polar molecule.

O<-N->CL

being polar more soluble in water4